OpenStudy (anonymous):

Difference Quotients Help. f(x)=1/x, f(x)-f(a)/x-a. I get how to plug in 1/x in the equation. Its the simplifying part that I stumble on.

OpenStudy (jamesj):

\[ \frac{ f(x) - f(a)}{x-a} = \frac{1/x - 1/a}{x-a} = \frac{(a-x)/(xa)}{x -a }\] Can you take it from there?

OpenStudy (anonymous):

what did you do between step 2 and 3?

OpenStudy (jamesj):

This: \[ \frac{1}{x} - \frac{1}{a} = \frac{a - x}{ax} \] That is, created a fraction with a common denominator

OpenStudy (anonymous):

can you show that step? did you take 1/x times a/a or what?

OpenStudy (jamesj):

You want a common denominator. That denominator has to be xa, because the denominators of 1/x and 1/a are x and a respectively. So how do you write 1/x as something/xa? It is 1/x = a/ax Likewise 1/a = x/ax Thus 1/x - 1/a = a/ax - x/ax Yes?

OpenStudy (anonymous):

ok that makes sense.

OpenStudy (anonymous):

Whats the easiest way to go from there?

OpenStudy (jamesj):

So we have shown that \[ \frac{f(x) - f(a)}{x-a} = \frac{(a-x)/xa}{x-a} = \frac{1}{xa} . \frac{a-x}{x-a} \] Looks like a pretty obviously cancellation to be done there!

OpenStudy (jamesj):

*obvious

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

The answer is -1/ax

OpenStudy (jamesj):

Yes